Carbon-14 has a radioactive half-life of 5730 years, which means every 5730 years half the Carbon-14 in a substance goes away. If a bone started with 100 mg of Carbon-14, but only has 10 mg left when it is discovered, how old is the bone?


1
Expert's answer
2020-09-28T08:05:43-0400

The following equation gives the quantitative relationship between the original number of nuclei present at time zero (N0) and the number (N) at a later time t:

"N = N_0e^{\u2212\u03bbt}"

λ is the decay constant

The relationship between the decay constant λ and the half-life t1/2 is

"λ = \frac{ln(2)}{t_{1/2}} ≈ \frac{0.693}{t_{1/2}}"

"λ = \frac{0.693}{5730} = 1.209\times10^{-4}"

"N\/N_0 = e^{\u2212\u03bbt}"

"10/100 = e^{−1.209\times10^{-4}t}"

"ln(0.1) = −1.209\times10^{-4}t"

"-2.302 = −1.209\times10^{-4}t"

"t = 1.904\times10^{4} = 19040\; years"